# Black Holes and Wormholes Beyond Classical General Relativity

### v1

@@ VI $F(R)$ modified gravitational models
Here, as done previously, we shall investigate approximate solutions, which are small deviations from the Scwharzschild solution.
Let us commence with the static metric as defined in Eq. (23). We shall deal again with a general case of modified gravity of the type $F(R)$ where Einstein-Hilbert Lagrangian is present, but in addition, higher-order curvature terms depending only on the Ricci scalar are added.
The simplest example is the well-known Starobinsky model for inflation [56] $F(R)=\gamma R+\omega R^{2}$, $\gamma=M^{2}_{P}$ Planck mass, and $\omega$ a dimensionless parameter.
@@ VI.1 Equations of Motion for Modified $F(R)$ Gravitational Models
Here, the following technique presented by Ref. [65] is valid for SSS. We will investigate mainly the Starobinsky model within SSSs, but the generalization to generic modified gravitational models is straightforward.
We introduce the quantity
@@
We have checked that these equations of motion reproduce the exact nontrivial BH solutions discussed in [66] within another approach. Recall that $F(R)=\gamma R+\omega R^{2}$, $\gamma=M^{2}_{P}$. When $\omega=0$, one has GR, with ${(F_{R})}_{0}=\gamma$ constant. The first equation gives $X_{0}=C_{0}$ constant. The second equation leads to
@@
We know that in GR, the unique solution is $f_{0}=1-\frac{r_{s}}{r}=g_{0}$. As a result, $C_{0}=X_{0}=1$. We can now look for approximate solutions for small $\omega$ of equations of motion (61a) and (61b). We make an expansion around the GR solution in the power series of the small coupling $\omega$ up to the second-order
$f=f_{0}+\omega f_{1}+\omega^{2}f_{2}+...\,,~{}X=1+\omega X_{1}+\omega^{2}X_{2}+...\,.$ (63)
@@
Up to the second order in $\omega$, the Eq. (61a) leads to
@@
from above equation and Eq. (60), one obtains
@@
At the second order in $\omega$, Eq. (61b) gives
@@
Now, let us discuss the existence of a horizon. One can obtained $r_{H}$ solving $g(r_{H})=0$, namely
@@
Thus, at this order, one again has a BH. In Appendix B, an alternative method is utilized to examine the modified gravitational model.
@@ VII Photon Rings and Shadows
In this Section, we will investigate the concept of photon ring and shadow associated with a generic SSS compact object, namely a BH or WH. The importance of this issue relies on the fact that strong gravity effects occur there. The topics are standard and can be found in several papers and textbooks (see, for example, [67, 68] and references therein). For the sake of completeness, we present, in the following, the main results.
@@
$\mathcal{L}=\frac{\dot{s}^{2}}{2}$
$=\frac{1}{2}\left(g_{tt}\dot{t}^{2}+2g_{t\rho}\dot{t}\dot{\rho}+g_{\rho\rho}\dot{\rho}^{2}+r(\rho)^{2}\left(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2}\right)\right)\,,$
@@
where we have introduced the impact parameter $b=\frac{L}{E}$.
For a mass-less particle ($k=0$), one gets
@@
These two equations permit the computation of $r_{c}$ and $b$ for mass-less particles. For example, for the Schwarzschild BH, one has $r=\rho$, $f=1-\frac{r_{s}}{r}$, $r_{c}=\frac{3r_{s}}{2}$, $b=\frac{3\sqrt{3}}{2}r_{s}$.
@@
WHs may be solutions of GR with exotic matter, but as discussed in the previous sections, solutions of modified gravity models or corrections to GR are also available. WHs admit photon spheres, which then possess shadows. For example, the Damour-Soludhukin WH has been investigated in [68].
As a first example of WH, let us consider now the $R^{2}$ static WH solution with $\rho=r$ and
@@
In the following, we compute the photon ring and the shadow of this compact object, its throat being $r_{H}=r_{s}$. The photon ring can be evaluated by making use of Eq. (91). Thus
@@
For $\lambda=0$, one recovers the GR result $r_{c}=\frac{3}{2}r_{s}$.
Assuming $\lambda$ is very small, one has
@@
$f(r)=1-\frac{2M}{r}+\omega f_{1}(r)\,,~{}g(r)=1-\frac{2M}{r}+\omega g_{1}(r)\,,\,$
$g_{1}(2M)=f_{1}(2M)\,.\quad\quad\quad\quad\quad\quad\quad\quad$ (100)
@@ VIII Conclusion
In this paper, first, we have revisited a non-singular static metric, including black holes not admitting a de Sitter core in the center and traversable wormhole, within a class of higher-order $F(R)$ modified gravity, satisfying the constraints $F(0)=\frac{dF}{dR}(0)=0$. These two conditions allow us to satisfy the equation of motion, even though a large class of $F(R)$ models is possible. Among them, the pure quadratic gravity model, namely $F(R)=R^{2}$, is of particular interest.
In the second part of the paper, making use of the so called effective field theory formulation of gravity, the quantum corrections to general relativity have been considered. We recall that within this approach, one starts with the Einstein-Hilbert action plus additive higher-derivative terms, typically built with curvature invariants with related arbitrary coupling constants. In our approach, such coupling constants have been considered small enough in order to justify linear and quadratic perturbation in the coupling constants around the GR solutions. In the important case of Static Spherically Symmetric space-times, it is well known that the zero-order solution is the Schwarzschild BH. In this paper, we have not investigated perturbation around the flat Minkowski solution.
In particular, in the case of Einstein-Hilbert action plus cubic curvature Goroff-Sagnotti contribution, the second-order corrections in the Goroff-Sagnotti coupling constant have been computed. In general, it has been shown that the effective metrics, namely Schwarzschild expression plus small quantum correction, are related to black holes and not to traversable wormholes. In this framework, within the approximation considered, the resolution of singularity for $r=0$ has not been resolved.
@@
Finally, with regard to future work, we think that one could try to go beyond the second-order approximation in the coupling constant in order to check if the black hole solutions are still present or wormhole solutions arise. The presence of wormhole solutions may occur since the higher-derivative terms can violate WEC (weak energy conditions), as shown in Sec. II and Appendix C. In fact, our formalism is suitable for this task, since, in the SSS case, we start with metric having the function $f(r)$ and $g(r)$ not proportional, namely with the choice $f(r)\neq g(r)b(r)$.
As far as the possible presence of WHs is concerned, we note that this fact has been numerically verified within Einstein-Weyl gravity (see, for example, [69] and references therein).
@@ Appendix A $\mathcal{C}^{n}$ theory
admits BH solution. The case $n=2$ has been investigated in Section IV.2.
At first order, one has
@@
$F_{n}(r)=-\frac{48^{n}(n-1)}{6(2n-1)}\left[1+2n+6(2n^{2}-n)\left(\frac{r}{M}-2\right)\right]\,,$ (105)
@@ Appendix B A specific class of modified gravitational models
$f(r)$ $=$ $f_{0}+\omega f_{1}+\omega^{2}f_{2}+..\,,$
$g(r)$ $=$ $g_{0}+\omega g_{1}+\omega^{2}g_{2}+..\,,$
$H$ $=$ $H_{0}+\omega H_{1}+\omega^{2}H_{2}+...\,.$ (112)
@@
$L$ $=$ $rg_{1}^{\prime}+g_{1}+H_{0}+\omega\frac{f_{1}-g_{1}}{2g_{0}}\left(rg_{1}^{\prime}+g_{1}+H_{0}\right)+$ (113)
$\omega(rg_{2}^{\prime}+g_{2}+H_{1})+\omega^{2}\left[\frac{f_{1}-g_{1}}{2g_{0}}\left(rg_{2}^{\prime}+g_{2}+H_{1}\right)\right]$
$+\omega^{2}\left[\frac{4f_{2}g_{0}-4g_{2}g_{0}-2f_{1}g_{1}-f_{1}^{2}+3g_{1}^{2}}{8g_{0}^{2}}\right]\times$
$\left(rg_{1}^{\prime}+g_{1}+H_{0}\right)+\omega^{2}\left(rg_{3}^{\prime}+g_{3}+H_{2}\right)+\mathcal{O}(\omega^{3})\,.$
@@
$L$ $=$ $rg_{1}^{\prime}+g_{1}+\omega\frac{f_{1}-g_{1}}{2g_{0}}\left(rg_{1}^{\prime}+g_{1}\right)+\omega(rg_{2}^{\prime}+g_{2})$ (114)
$+\omega^{2}\frac{f_{1}-g_{1}}{2g_{0}}\left(rg_{2}^{\prime}+g_{2}\right)$
$+\omega^{2}\left[\frac{4f_{2}g_{0}-4g_{2}g_{0}-2g_{1}f_{1}-f_{1}^{2}+3g_{1}^{2}}{8g_{0}^{2}}\right]\times$
$\left(rg_{1}^{\prime}+g_{1}\right)+\omega^{2}(rg_{3}^{\prime}+g_{3}+H_{2})+\mathcal{O}(\omega^{3})\,.$
@@
$H$ $=$ $r^{2}(R_{0}^{2}+2\omega R_{0}R_{1}+\omega^{2}R_{2})\,,\quad H_{0}=r^{2}R^{2}_{0}\,,$
$H_{1}$ $=$ $2r^{2}R_{0}R_{1}\,,\quad H_{2}=r^{2}R_{1}^{2}\,.$ (115)
@@
$H=r^{2}f(R_{0}+\omega R_{1})=r^{2}f(R_{0})+r^{2}\omega R_{1}f^{\prime}(R_{0})$
$+r^{2}\omega^{2}\frac{f^{\prime\prime}(R_{0})}{2}R_{1}^{2}+...\,.$ (117)
Thus $H_{0}=r^{2}f(R_{0})\,,~{}H_{1}=r^{2}f^{\prime}(R_{0})R_{1}\,,~{}H_{2}=r^{2}\frac{f^{\prime\prime}(R_{0})}{2}R_{1}^{2}$. In the Schwarzschild case, $R_{0}=0$, and $H_{0}=H_{1}=0$.
@@
First, we may redefine the time in order to deal with an asymptotic flat Minkowski space-time, $dT^{2}=(1+\omega c_{2})dt^{2}\,c_{2}>0,~{}\omega>0$. Thus, we arrive at
@@
Which has to be interpreted as vanishing at first order in $\omega$. Thus one concludes that we do not have a generalized Damour-Solodhukin WH [70] as soon as $c_{1}<0$, but one is dealing with a Schwarzschild-like BH. In fact in the linear $\omega$ approximation, we have
@@
Thus, we are left with $f_{1}=g_{1}=\frac{c_{1}}{r}$.
At the second order in $\omega$, one has
$rg_{2}^{\prime}+g_{2}=4g_{0}\bigg{(}-r^{2}R_{1}\frac{\partial{R_{1}}}{\partial f_{1}}+\frac{d}{dr}\left(r^{2}R_{1}\frac{\partial{R_{1}}}{\partial f_{1}^{\prime}}\right)$
$-\frac{d^{2}}{dr^{2}}\left(r^{2}R_{1}\frac{\partial{R_{1}}}{\partial f_{1}^{\prime\prime}}\right)\bigg{)}=G_{2}(r)\,.\quad$ (127)
@@
$\frac{d}{dr}\left(\frac{r(f_{2}-g_{2})}{g_{0}}\right)-\frac{f_{2}-g_{2}}{g_{0}}=-G_{2}(r)$
$+r^{2}\frac{\partial R_{1}}{\partial g_{1}}-\frac{d}{dr}\left(r^{2}\frac{\partial R_{1}}{\partial g_{1}^{\prime}}\right)\,.$ (129)
@@ Appendix C Buchdahl wormhole solution
Now we consider the non-trivial spherically symmetric static solution of pure $R^{2}$ gravity found in [71, 72]. In isotropic coordinates, one has
$F(\rho)$ $=$ $\left(\frac{4\rho-ar_{s}}{4\rho+ar_{s}}\right)^{\alpha}\,,$
$G(\rho)$ $=$ $\left(\frac{1}{4\rho}\right)^{4}(4\rho+ar_{s})^{2-\beta}(4\rho-ar_{s})^{2+\beta}\,,$ (130)
@@
Above, $b$ is a dimensionless free parameter (the Buchdahl parameter), and $r_{s}$ is the constant radius. When the Buchdahl parameter is vanishing, $a=1$, $\alpha=2$, $\beta=-2$ and one recovers the Schwarzschild solution (13). The areal radius reads
@@
The location of the horizon is given by solving the Eq. (10), or alternatively $(\partial_{\rho}r)^{2}=0$. Thus, one has

### v2

@@ VI $F(R)$ modified gravitational models
Here, as done previously, we shall investigate approximate solutions, which are small deviations from the Scwharzschild solution. We shall commence with the static metric as defined in equation (27). We shall deal again with a general case of modified gravity of the type $F(R)$ where Einstein-Hilbert Lagrangian is present, but in addition, higher-order curvature terms depending only on the Ricci scalar are added. The simplest example is the well-known Starobinsky model for inflation [58] $F(R)=\gamma R+\omega R^{2}$, $\gamma=M^{2}_{P}$ Planck mass, and $\omega$ a dimensionless parameter.
@@ VI.1 Equations of Motion for Modified $F(R)$ Gravitational Models
Here, the following technique presented by the Ref. [67] is valid for SSS. We will investigate mainly the Starobinsky model within SSSs, but the generalization to generic modified gravitational models is straightforward. We introduce the quantity
@@
We have checked that these equations of motion reproduce the exact nontrivial BH solutions discussed in [44] within another approach. Recall that $F(R)=\gamma R+\omega R^{2}$, $\gamma=M^{2}_{P}$ When $\omega=0$, one has GR, with ${(F_{R})}_{0}=\gamma$ constant. The first equation gives $X_{0}=C_{0}$ constant. The second equation leads to
@@
We know that in GR, the unique solution is $f_{0}=1-\frac{r_{s}}{r}=g_{0}$. As a result, $C_{0}=X_{0}=1$. We can now look for approximate solutions for small $\omega$ of equations of motion (65a) and (65b). An expansion around the GR solution in the power series of the small coupling $\omega$ up to the second-order
$f=f_{0}+\omega f_{1}+\omega^{2}f_{2}+...\,,\quad X=1+\omega X_{1}+\omega^{2}X_{2}+...\,.$ (67)
@@
Up to the second order in $\omega$, the equation (65a) leads to
@@
from above equation and equation (64), one obtains
@@
At the second order in $\omega$, equation (65b) gives
@@
Now, we shall discuss the existence of a horizon. One can obtained $r_{H}$ solving $g(r_{H})=0$, namely
@@
Thus, at this order, one again has a BH.
@@ VII Photon Rings and Shadows
In this section, we will investigate the concept of photon ring and shadow associated with a generic SSS compact object, namely a BH or WH. The important of this issue relies on the fact that strong gravity effects occur there. The topics are standard and can be found in several papers and textbooks (see, for example, [68, 69] and references therein). For the sake of completeness, we present, in the following, the main results.
@@
$\mathcal{L}=\frac{\dot{s}^{2}}{2}=\frac{1}{2}\left(g_{tt}\dot{t}^{2}+2g_{t\rho}\dot{t}\dot{\rho}+g_{\rho\rho}\dot{\rho}^{2}+r(\rho)^{2}\left(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2}\right)\right)\,,$ (85)
@@
where we have introduced the impact parameter $b=\frac{L}{E}$. For a mass-less particle ($k=0$), one gets
@@
These two equations permit the computation of $r_{c}$ and $b$ for mass-less particles. For example, for the Scwarschild BH, one has $r=\rho$, $f=1-\frac{r_{s}}{r}$, $r_{c}=\frac{3r_{s}}{2}$, $b=\frac{3\sqrt{3}}{2}r_{s}$.
@@
WHs may be solutions of GR with exotic matter, but as discussed in the previous sections, solutions of modified gravity models or corrections to GR are also available. WHs admit photon spheres, which then possess shadows. For example, the Damour-Soludhukin WH has been investigated in [69]. As a first example of WH, let us consider now the $R^{2}$ static WH solution with $\rho=r$ and
@@
In the following, we compute the photon ring and the shadow of this compact object, its throat being $r_{H}=r_{s}$. The photon ring can be evaluated by making use of equation (95). Thus
@@
For $\lambda=0$, one recovers the GR result $r_{c}=\frac{3}{2}r_{s}$. Assuming $\lambda$ is very small, one has
@@
$f(r)=1-\frac{2M}{r}+\omega f_{1}(r)\,,\quad g(r)=1-\frac{2M}{r}+\omega g_{1}(r)\,,\,\quad g_{1}(2M)=f_{1}(2M)\,.$ (104)
@@
The recent papers have been published, providing bounds on the model parameters $M$. For further details, please refer to the papers by [73, 72].
@@ VIII Conclusion
In this paper, first we have revisited a non-singular static metric, including black holes not admitting a de Sitter core in the centre and traversable wormhole, within a class of higher-order $F(R)$ modified gravity, satisfying the constraints $F(0)=\frac{dF}{dR}(0)=0$. These two conditions allow to satisfy the equation of motions, even though a large class of $F(R)$ models are possible. Among them, the pure quadratic gravity model, namely $F(R)=R^{2}$ is of particular interest.
In the second part of the paper, making use of the so called effective field theory formulation of gravity, the quantum corrections to GR has been considered. We recall that within this approach one starts with the Einstein-Hilbert action plus additive higher-derivative terms, typically built with curvature invariants, with related arbitrary coupling constants. In our approach, such coupling constants have been considered small enough in order to justify linear and quadratic perturbation in the coupling constants around the GR solutions. In the important case of SSSs, it is well known that the zero order solution is the Schwarzschild BH. In this paper, we have not investigated perturbation around the flat Minkoswki solution.
In particular, in the case of Einstein-Hilbert action plus cubic curvature Goroff-Sagnotti contribution, the second order corrections in the Goroff-Sagnotti coupling constant have been computed. In general, it has been shown that the effective metrics, namely Schwarschild expression plus small quantum correction are related to black holes, and not to traversable wormholes. In this framework, within the approximation considered, the resolution of singularity for $r=0$ has not been resolved.
@@
Finally, with regard to future work, we think that one could try to go beyond the second order approximation in the coupling constant, in order to check if the black hole solutions are still present, or wormholes solutions arise. The presence of wormholes solution may occur, since the higher-derivative terms can violate WEC (weak energy conditions), as shown in Section II and appendix C. In fact, our formalism is suitable for this task, since, in the SSS case, we start with metric having the function $f(r)$ and $g(r)$ not proportional, namely with the choice $f(r)\neq g(r)b(r)$.
As far as the possible presence of WHs is concerned, we note that this fact has been numerically verified within Einstein-Weyl gravity (see for example [74] and references therein) .
@@ Appendix A $\mathcal{C}^{n}$ theory
admits BH solution. The case $n=2$ has been investigated in Section IV.2. At first order, one has
@@
$F_{n}(r)=-\frac{48^{n}(n-1)}{6(2n-1)}\left(1+2n+(12n^{2}-6n)\left(\frac{r}{M}-2\right)\right)\,,$ (109)
@@ Appendix B A specific class of modified gravitational models
$f(r)=f_{0}+\omega f_{1}+\omega^{2}f_{2}+..\,,~{}~{}~{}g(r)=g_{0}+\omega g_{1}+\omega^{2}g_{2}+..\,,~{}~{}~{}H=H_{0}+\omega H_{1}+\omega^{2}H_{2}+...\,.$ (116)
@@
$L$ $=$ $rg_{1}^{\prime}+g_{1}+H_{0}+\omega\frac{f_{1}-g_{1}}{2g_{0}}\left(rg_{1}^{\prime}+g_{1}+H_{0}\right)+\omega(rg_{2}^{\prime}+g_{2}+H_{1})+\omega^{2}\left(\frac{f_{1}-g_{1}}{2g_{0}}\left(rg_{2}^{\prime}+g_{2}+H_{1}\right)\right)$ (117)
$+\omega^{2}\left(\frac{4f_{2}g_{0}-4g_{2}g_{0}-2f_{1}g_{1}-f_{1}^{2}+3g_{1}^{2}}{8g_{0}^{2}}\right)\left(rg_{1}^{\prime}+g_{1}+H_{0}\right)+\omega^{2}\left(rg_{3}^{\prime}+g_{3}+H_{2}\right)+\mathcal{O}(\omega^{3})\,.$
@@
$L$ $=$ $rg_{1}^{\prime}+g_{1}+\omega\frac{f_{1}-g_{1}}{2g_{0}}\left(rg_{1}^{\prime}+g_{1}\right)+\omega(rg_{2}^{\prime}+g_{2})+\omega^{2}\frac{f_{1}-g_{1}}{2g_{0}}\left(rg_{2}^{\prime}+g_{2}\right)$ (118)
$+\omega^{2}\left(\frac{4f_{2}g_{0}-4g_{2}g_{0}-2g_{1}f_{1}-f_{1}^{2}+3g_{1}^{2}}{8g_{0}^{2}}\right)\left(rg_{1}^{\prime}+g_{1}\right)+\omega^{2}(rg_{3}^{\prime}+g_{3}+H_{2})+\mathcal{O}(\omega^{3})\,.$
@@
$H=r^{2}(R_{0}^{2}+2\omega R_{0}R_{1}+\omega^{2}R_{2})\,,\quad H_{0}=r^{2}R^{2}_{0}\,,\quad H_{1}=2r^{2}R_{0}R_{1}\,,\quad H_{2}=r^{2}R_{1}^{2}\,.$ (119)
@@
$H=r^{2}f(R_{0}+\omega R_{1})=r^{2}f(R_{0})+r^{2}\omega R_{1}f^{\prime}(R_{0})+r^{2}\omega^{2}\frac{f^{\prime\prime}(R_{0})}{2}R_{1}^{2}+...\,.$ (121)
Thus $H_{0}=r^{2}f(R_{0})\,,~{}H_{1}=r^{2}f^{\prime}(R_{0})R_{1}\,,~{}H_{2}=r^{2}\frac{f^{\prime\prime}(R_{0})}{2}R_{1}^{2}$. In the Scwharzschild case, $R_{0}=0$, and $H_{0}=H_{1}=0$.
@@
First, we may redefine the time in order to deal with an asymptotic flat Minkowski space-time, $dT^{2}=(1+\omega c_{2})dt^{2}\,c_{2}>0,~{}\omega>0$. Thus we arrive at
@@
Which has to be interpreted as vanishing at first order in $\omega$. Thus one concludes that we do not have a generalized Damour-Solodhukin WH [70] as soon as $c_{1}<0$, but one is dealing with a Schwarschild-like BH. In fact in the linear $\omega$ approximation, we have
@@
Thus, we are left with $f_{1}=g_{1}=\frac{c_{1}}{r}$. At the second order in $\omega$, one has
$rg_{2}^{\prime}+g_{2}=4g_{0}\left(-r^{2}R_{1}\frac{\partial{R_{1}}}{\partial f_{1}}+\frac{d}{dr}\left(r^{2}R_{1}\frac{\partial{R_{1}}}{\partial f_{1}^{\prime}}\right)-\frac{d^{2}}{dr^{2}}\left(r^{2}R_{1}\frac{\partial{R_{1}}}{\partial f_{1}^{\prime\prime}}\right)\right)=G_{2}(r)\,.$ (131)
@@
$\frac{d}{dr}\left(\frac{r(f_{2}-g_{2})}{g_{0}}\right)-\frac{f_{2}-g_{2}}{g_{0}}=-G_{2}(r)+r^{2}\frac{\partial R_{1}}{\partial g_{1}}-\frac{d}{dr}\left(r^{2}\frac{\partial R_{1}}{\partial g_{1}^{\prime}}\right)\,.$ (133)
@@ Appendix C Buchdahl wormhole solution
Now we consider the non-trivial SSS solution of pure $R^{2}$ gravity found in [71]. In isotropic coordinates, one has
$F(\rho)=\left(\frac{4\rho-ar_{s}}{4\rho+ar_{s}}\right)^{\alpha}\,,\quad G(\rho)=\left(\frac{1}{4\rho}\right)^{4}(4\rho+ar_{s})^{2-\beta}(4\rho-ar_{s})^{2+\beta}\,,$ (134)
@@
Above, $b$ is a dimensionless free parameter (the Buchdahl parameter), and $r_{s}$ is the constant radius. When the Buchdahl parameter is vanishing, $a=1$, $\alpha=2$, $\beta=-2$ and one recovers the Scwarzschild solution (17). The areal radius reads
@@
The location of the horizon is given by solving the equation (14), or alternatively $(\partial_{\rho}r)^{2}=0$. Thus, one has